solve and check
6(2 – x) = 4x

Answer by stanbon(75887)

solve and check
6(2 – x) = 4x
12 - 6x = 4x
12 = 10x
x= 6/5
================
Checking:
6(2-(6/5)) = 4(6/5)
12-36/5 = 24/5
60/5 - 36/5 = 24/5
24/5 = 24/5
=================
Cheers,
Stan H.